# DINAMO: Dynamic and INterpretable Anomaly MOnitoring for Large-Scale Particle Physics Experiments

### v1

@@ Abstract
Ensuring reliable data collection in large-scale particle physics experiments demands Data Quality Monitoring (DQM) procedures to detect possible detector malfunctions and preserve data integrity. Traditionally, this resource-intensive task has been handled by human shifters that struggle with frequent changes in operational conditions. We present novel, interpretable, robust, and scalable DQM algorithms designed to automate anomaly detection in time-dependent settings. Our approach constructs evolving histogram templates with built-in uncertainties, featuring both a statistical variant — extending the classical Exponentially Weighted Moving Average (EWMA) — and a machine learning (ML)-enhanced version that leverages a transformer encoder for improved adaptability. Experimental validations on synthetic datasets demonstrate the high accuracy, adaptability, and interpretability of these methods, with the statistical variant being commissioned in the LHCb experiment at the Large Hadron Collider, underscoring its real-world impact. The code used in this study is available at https://github.com/ArseniiGav/DINAMO.
@@
Anomaly Detection; Interpretability; Online Learning; EWMA; Transformer; High Energy Physics, Data Quality Monitoring; Robustness
I Introduction
Large-scale particle physics experiments, particularly at the Large Hadron Collider (LHC) [1, 2, 3, 4], collect immense amounts of data that must be continuously validated to ensure its integrity. Traditionally, this Data Quality Monitoring (DQM) task has relied on dedicated human shifters, who review histograms and other summary statistics to spot anomalous behavior in detector outputs. Such manual monitoring is both resource-intensive—requiring large teams of non-expert personnel over extended periods—and prone to human error, fatigue, and inconsistency.
DQM generally takes place in two main stages: an online stage, which runs in near real-time during data collection to quickly flag issues, and an offline stage, which offers a more exhaustive evaluation after data-taking has concluded. Although automated approaches have been proposed, a persistent challenge is that the so-called “nominal” detector behavior is not static: operational conditions can shift significantly over time, especially after substantial hardware or software interventions. Indeed, the LHC experiments are undergoing major upgrades in preparation for the High-Luminosity LHC era [5], and the commissioning periods that follow such upgrades often entail frequent, sometimes unforeseen, modifications to detector configurations. In that scenario, a successful automated DQM solution must satisfy several stringent requirements:
  • High Accuracy: It should reliably distinguish between good and bad data. Discarding valid data reduces valuable physics potential, whereas accepting faulty data risks misleading scientific conclusions.
  • Targeted Human Oversight: While human expertise is indispensable for diagnosing complex issues, it should be engaged only when necessary. Data that agree with nominal references (within uncertainties) should not burden shifters.
  • Adaptability to Changing Conditions: The reference model for nominal data should be updated seamlessly as operational settings evolve. Manual updating is labor-intensive and prone to error.
  • Interpretability: Humans must be able to understand and verify decisions made by automated algorithms. In addition, providing supplementary information that helps humans pinpoint the root causes of flagged anomalies is highly valuable.
In this paper, we introduce DINAMO (Dynamic and INterpretable Anomaly MOnitoring), a framework that learns statistical representations of nominal data in histogram format. By maintaining templates and explicit uncertainties, we enable natural fluctuations or benign variations (e.g., detector noise) to be distinguished from genuine anomalies. Once constructed, these templates serve as references for newly observed histograms, triggering an alert if the data deviate beyond expected bounds. Importantly, the templates themselves provide an interpretable visualization for domain experts to verify and debug the system’s outputs.
We present two algorithmic variants:
  • DINAMO-S: A statistics-based approach that leverages new generalizations of Exponentially Weighted Moving Averages (EWMA). It offers simplicity, scalability, and ease of deployment, making it well-suited to rapid adoption in particle physics experiments.
  • DINAMO-ML: An ML-enhanced architecture employing a Transformer encoder, which adaptively learns to combine past good histograms and track more nuanced temporal or bin-level dependencies. While more resource-intensive than DINAMO-S, it provides higher accuracy and faster adaptation.
We validate these algorithms on a suite of synthetic datasets designed to mimic both gradual drifts and sudden operational shifts, alongside a diverse spectrum of anomaly types. The datasets capture the main characteristics of real detector data in a generalizable fashion. Notably, DINAMO-S is being commissioned in the LHCb collaboration [3] at the Large Hadron Collider for offline DQM in an upcoming data-taking period, underscoring its practical relevance and real-world effectiveness.
This paper is organized as follows: Section II reviews related work in anomaly detection and adaptive monitoring, both within experimental particle physics and in general. Section III describes the statistical and ML-based methods, including their mathematical formulation and design principles. Section IV presents the experimental setup used to study the behavior of the presented algorithms, including the synthetic dataset design, evaluation protocols, and results. Section V discusses the findings, addresses limitations, and elaborates on the application at the LHCb experiment. Finally, Section VI summarises the conclusions.
II Related Work
Anomaly detection in particle physics experiments often relies on reference-based statistical tests (e.g. $\chi^{2}$, Kolmogorov–Smirnov) against “golden” or pre-certified datasets. Increasingly, machine-learning approaches based on supervised CNN classifiers, autoencoders, and LSTMs have also been employed for DQM tasks [6]. A recent example is the usage of autoencoders to flag anomalous calorimeter readouts at the CMS Collaboration [7, 2]. Other recent strategies, like “AutoDQM” [8], demonstrate the power of combining advanced statistical techniques with unsupervised machine-learning methods. However, these approaches typically assume relatively stable operational regimes and sporadic reference updates, which can break down in the commissioning phases of new of upgraded detectors, where unforeseen hardware/software changes occur frequently. Reinforcement learning (RL) approaches [9] have recently been proposed to handle time-varying conditions and further automate the DQM process, though these pipelines can introduce significant complexity and may require substantial time for experiments to adopt.
@@ II Related Work
Beyond particle physics, many time series anomaly-detection frameworks track evolving data distributions via classical methods like the Exponentially Weighted Moving Average (EWMA) [10], which can flag deviations from a running mean but may struggle with nonlinear drifts or high-dimensional data. To model more complex behaviors, machine-learning architectures—ranging from RNNs and LSTMs [11, 12] to Transformers [13, 14]—have become prominent, although they often require frequent retraining or threshold tuning when distributions shift. Online or incremental learning [15] offers an option for parameters to adapt continuously to new data in dynamic settings. Finally, a persistent challenge lies in the interpretability of the tools, as many contemporary algorithms produce opaque anomaly scores [16]. In safety-critical domains like experimental particle physics, explainable anomaly detection [17]—with explicit uncertainties or feature-level deviation maps—remains essential for building trust and enabling root-cause analysis.
III Methods
In this section, we describe the proposed DINAMO-S and DINAMO-ML algorithms. We begin by outlining the system’s core notation and assumptions, then present a high-level description of each algorithm, including their update rules and anomaly detection strategies. Complete mathematical details, full pseudocode, and additional technical considerations are deferred to the appendices.
III.1 System Overview and Notation
We consider a sequence of runs indexed by $i=1,2,\ldots,n$. In the context of particle physics experiments, a run corresponds to a data-collection interval during which detectors accumulate events. These events can be integrated and visualized via histograms that capture key detector responses (e.g., energy deposits, track multiplicities, etc.) over the course of that run. Runs are ordered sequentially, following the data-collection order. Particle physics experiments typically present two levels of monitoring:
  • Online Monitoring: Performed in real time while data is being collected. The main goal is to identify anomalies as quickly as possible, in order to fix the underlying cause. Given the time constraints, this monitoring is not as exhaustive as the one offline.
  • Offline Monitoring: Carried out after the data has been recorded, allowing for a more thorough search for anomalies. Only data passing this second check is used in subsequent scientific studies.
@@ III.1 System Overview and Notation
Our proposed algorithm is designed to be applicable in either regime. It assumes an operational setup where a trained model quickly produces an anomaly score for each new run. If the score is under a certain threshold, the data can be automatically accepted. Otherwise, the system escalates the run to human experts, displaying its prediction along with summary plots or statistics. The human shifter can then decide whether the run is genuinely anomalous, or if the experimental conditions have simply changed in a way that the algorithm has yet to incorporate into its reference model.
@@
Formally, we denote each one-dimensional histogram for run $i$ by $\bm{x}_{i}\in\mathbb{R}^{N_{b}}$, where $N_{b}$ is the number of bins. We associate each run with a label $y_{i}\in\{0,1\}$, indicating whether the histogram is deemed good (0) or bad (1). This label may come from a human shifter (in real deployments) or from ground truth (in synthetic experiments). A good label typically implies that the detector output conforms to nominal expectations, whereas a bad label signifies that significant anomalies or faulty conditions are present. Good histograms can vary from run to run due to per-bin statistical and intrinsic uncertainties. In this paper, the concept of intrinsic uncertainty is meant to cover any type of fluctuation in detector response that is not deemed to impact scientific analysis and is hence harmless.
A central concept in our framework is the reference histogram $\bm{\mu}$ and its per-bin uncertainties $\bm{\sigma_{\mu}}$, which represent the current best estimate of the nominal (i.e., good) distribution for a given type of histogram. When a newly observed histogram $\bm{x}_{i}$ is confirmed to be good, the reference is updated to incorporate this information; if a histogram is deemed bad, the reference remains unchanged. Both DINAMO-S and DINAMO-ML produce a $\chi^{2}$-based anomaly score that helps decide if a histogram is consistent with the reference.
III.2 DINAMO-S: Statistical Method

### v2

@@ Abstract
Ensuring reliable data collection in large-scale particle physics experiments demands Data Quality Monitoring (DQM) procedures to detect possible detector malfunctions and preserve data integrity. Traditionally, this resource-intensive task has been handled by human shifters who struggle with frequent changes in operational conditions. We present DINAMO: a novel, interpretable, robust, and scalable DQM framework designed to automate anomaly detection in time-dependent settings. Our approach constructs evolving histogram templates with built-in uncertainties, featuring both a statistical variant — extending the classical Exponentially Weighted Moving Average (EWMA) — and a machine learning (ML)-enhanced version that leverages a transformer encoder for improved adaptability. Experimental validations on synthetic datasets demonstrate the high accuracy, adaptability, and interpretability of these methods. The statistical variant is being commissioned in the LHCb experiment at the Large Hadron Collider, underscoring its real-world impact. The code used in this study is available at https://github.com/ArseniiGav/DINAMO.
@@
Keywords: anomaly detection, interpretability, online learning, EWMA, transformer, high energy physics, data quality monitoring
1 Introduction
The integrity of experimental data is fundamental to the success of any scientific discovery in physics. In large-scale particle physics experiments, particularly at the Large Hadron Collider (LHC) [1, 2, 3, 4], vast amounts of data are collected continuously, requiring robust validation mechanisms to ensure their quality. The complex, multicomponent nature of such experimental instruments makes various types of failures at different levels of data collection inevitable. To ensure that the physics experiments bring reliable scientific discoveries, strict quality criteria are applied to all collected data, with only validated data proceeding to subsequent physics analyses. This monitoring and data certification process is done through data quality monitoring (DQM) systems [5]. Traditionally, DQM systems rely on dedicated human shifters who review aggregated information from specific detector components or variables to identify anomalous behavior in detector outputs. Following specific predetermined guidelines, the identification is then performed by comparing collected data to the expected (so-called reference) detector outputs. The references are typically constructed manually by human experts based on previously validated good data. This manual approach presents several challenges:
  • Firstly, it requires significant human resources. The complexity and volume of data often exceed what human operators can efficiently process leading to potential errors and inconsistencies. Moreover, interpretation differences between shifters can result in inconsistent and untraceable quality assessments.
  • Secondly, while human shifters are often not detector experts, reference construction and guidelines formulation require consistent attention from a small group of detector experts. This is particularly problematic during dynamic operational periods, such as the commissioning of a new detector component or a major software update.
DQM typically operates on two levels: an online stage, which runs in near real-time during data collection to quickly flag issues [6, 7], and an offline stage, which offers a more thorough evaluation after a data-taking run has concluded [8]. Both stages are similar but differ in a few important aspects: the online stage requires immediate shifter response and operates under computational constraints that limit the scope of monitored features, whereas the offline stage allows comprehensive analysis without the direct time pressure. Usually, the final certification of the collected data is done at the offline stage. At both levels, an effective automated DQM solution must satisfy several key requirements:
  • High Accuracy: The system must reliably distinguish between good and bad data. False positives waste valuable physics data, while false negatives risk compromising scientific conclusions.
  • Targeted Human Oversight: While human expertise remains essential for complex diagnostics, it should be engaged selectively. Data statistically consistent with nominal references can be automatically approved, freeing human resources for more critical tasks.
  • Adaptability: References and their uncertainties should ideally update seamlessly as operational conditions evolve, without requiring manual intervention.
  • Interpretability: The system’s decisions must be transparent and verifiable by detector experts. Algorithms should provide valuable information to help identify anomaly causes and enable verification of decision-making processes.
Data quality monitoring intersects with broader anomaly detection research. In particle physics, traditional approaches rely on reference-based statistical tests (e.g., $\chi^{2}$, Kolmogorov–Smirnov) against pre-certified “golden” datasets. More recently, machine learning techniques have gained prominence, including supervised CNN classifiers, autoencoders, and LSTMs [5]. The CMS Collaboration [2], for instance, has employed autoencoders to identify anomalous calorimeter readouts [9], while a novel approach called “AutoDQM” [10] combines statistical techniques with unsupervised machine learning.
However, existing automated approaches typically assume static operational regimes, with ML-based methods often requiring re-training when conditions change. While effective for relatively stable conditions, these methods become less pragmatic during detector commissioning periods characterized by frequent and unpredictable operational changes. This limitation is particularly evident during commissioning periods following major upgrades, such as those underway for the High-Luminosity LHC era [11]. These periods are characterized by frequent, sometimes unexpected, modifications to detector configurations. While reinforcement learning methods have been proposed to handle time-dependency [12] and further automate the DQM process, their complexity and the substantial time required for experiments can be barriers to adoption.
@@ 1 Introduction
Beyond particle physics, many time series anomaly-detection frameworks track evolving data distributions via classical methods like the Exponentially Weighted Moving Average (EWMA) [13], which can flag deviations from a running mean but may struggle with nonlinear drifts or high-dimensional data. To model more complex behaviors, machine-learning architectures — ranging from RNNs and LSTMs [14, 15] to Transformers [16, 17] — have become prominent, although they often require frequent re-training or threshold tuning when distributions shift. Online or incremental learning [18] offers an option for parameters to adapt continuously to new data in dynamic settings. Finally, a persistent challenge lies in the interpretability of the tools, as many contemporary algorithms produce opaque anomaly scores [19]. In safety-critical domains like experimental particle physics, explainable anomaly detection [20] — with explicit uncertainties or feature-level deviation maps—remains essential for building trust and enabling root-cause analysis.
In this paper, we introduce DINAMO (Dynamic and INterpretable Anomaly MOnitoring), a framework comprising two complementary approaches for automating DQM in particle physics. Our solution emphasizes adaptability to evolving detector conditions and interpretability while maintaining robustness and technical simplicity. The framework includes a statistical method based on a modification of EWMA (DINAMO-S) and a machine learning (ML) method that substitutes the EWMA with a transformer encoder (DINAMO-ML). Although developed with the LHCb experiment [3] as the primary use case, DINAMO is designed to be general enough for deployment across different experiments in both online and offline regimes. Demonstrating its practical impact in operational settings, DINAMO-S is currently being commissioned at the LHCb experiment for offline DQM. To enable comprehensive evaluation of the methods, we developed a synthetic data generator that reproduces gradual drifts, sudden operational shifts, systematic uncertainties, and other features of real-world detector data.
The rest of this paper is organized as follows: Section 2 presents details of the DINAMO approaches, their technical aspects and mathematical formalization. Section 3 presents the evaluation metrics. Section 4 describes the synthetic data generator and its features. Section 5 is dedicated to the results and to the methods comparison. Section 6 presents the summary, discusses the limitations of the methods and the practical implementation at LHCb.
2 Methods
In this section, we describe the proposed DINAMO-S and DINAMO-ML algorithms. DINAMO learns statistical representations of nominal data in histogram format, maintaining templates with explicit uncertainties that distinguish between statistical fluctuations and genuine anomalies. These templates serve as references for newly observed histograms, triggering alerts when data is not statistically compatible (accounting for the uncertainties). Furthermore, the uncertainties of the templates serve a dual purpose: they capture acceptable variations in detector response (such as degenerate good solutions where noisy pixels induce visible but acceptable deviations), and when combined with test statistics, they provide the foundation for direct interpretability — enabling domain experts to verify system outputs and perform root-cause analysis.
In the context of particle physics experiments, a run corresponds to a data-collection interval during which detectors accumulate events. Each run is monitored through a collection of histograms, where the bin counts accumulate information over the whole duration of the run. Therefore, different run durations result in varying total counts and so varying bin-to-bin statistical uncertainties. While we model these uncertainties using Poisson statistics in this work, the DINAMO framework is generalizable to histograms with any type of per-bin uncertainties.
As discussed in Section 1, the framework comprises two complementary approaches:
  • DINAMO-S: A non-ML, statistics-based approach that extends EWMA with additional statistical re-weighting factors to account for the Poisson uncertainty of good runs. It offers simplicity, scalability, and ease of deployment, potentially enabling rapid adoption in experimental workflows.
  • DINAMO-ML: An ML-based approach that uses a deep neural network that is trained via online learning to predict future references and their uncertainties while taking as input past good histograms. Computationally more intensive than DINAMO-S, it is designed to offer higher accuracy and faster adaptation to changing conditions.
@@ 2 Methods
We consider a sequence of runs indexed by $i=1,2,\ldots,n$. These events can be integrated and visualized via histograms that capture key detector responses (e.g., energy deposits, track multiplicities, etc.) over the course of that run. Runs are ordered sequentially, following the data-collection order. Our proposed algorithms are designed to be applicable in both online and offline regimes of DQM. They assume an operational setup where a trained model quickly produces an anomaly score for each new run. If the score is under a certain threshold, the data can be automatically accepted; otherwise, the system can send the run to human experts, displaying its prediction along with summary plots or statistics. The human shifter can then decide whether the run is genuinely anomalous, or if experimental conditions have changed in a way the algorithm has not yet learned.
@@
Formally, we denote the one-dimensional histogram for run $i$ by $\bm{x}_{i}\in\mathbb{R}^{N_{b}}$, where $N_{b}$ is the number of bins. We associate each run with a label $y_{i}\in\{0,1\}$, indicating whether the histogram is good (0) or bad (1). This label may come from a human shifter (in real deployments) or from ground truth (in synthetic experiments). A good label typically implies that the detector output conforms to nominal expectations, whereas a bad label signifies that significant anomalies or faulty conditions are present. Good histograms can vary from run to run due to per-bin statistical and intrinsic uncertainties. Here, the concept of intrinsic uncertainty is meant to cover any type of fluctuation in detector response that is not impacting the subsequent analysis and is hence harmless.
A central concept in our framework is the reference histogram $\bm{\mu}$ and its per-bin uncertainties $\bm{\sigma_{\mu}}$, which represent the current best estimate of the nominal (i.e., good) distribution for a given type of histogram. When a newly observed histogram $\bm{x}_{i}$ is confirmed to be good, the reference is updated to incorporate this information; if a histogram is bad, the reference remains unchanged. Both DINAMO-S and DINAMO-ML produce a $\chi^{2}$-based anomaly score that helps decide if a histogram is consistent with the reference.
2.1 DINAMO-S: Statistical Method